Sao is 4 times older than Agnes at present. Ten years ago,Sao was 12 times older than her.How old are Sao and Agnes now?

Answer by sachi(548)

Sao is 4 times older than Agnes at present. Ten years ago,Sao was 12 times older than her.How old are Sao and Agnes now?

ans:
let Agnes is x yrs old at present
so Sao is 4x yrs old now
ten yrs ago
Agnes was x-10 yrs old
& Sao was 4x-10 yrs old
as per question 4x-10=12(x-10)=12x-120
or 12x-4x=120-10
or 8x=110
or x=110/8=13.75 yrs
& y=4*13.75=55 yrs ans